OpenStudy (anonymous):

How could it be that Sum(n^2/(n^3+1)) from n=0 to infinity is divergent?

OpenStudy (anonymous):

For clarity: \[ \sum\limits_{n=0}^\infty \frac{n^2}{n^3 + 1} \]

OpenStudy (anonymous):

Wait sorry, completely forgot. Of course that series shouldn't converge! But it was alternating: \[ \sum\limits_{n=0}^\infty (-1)^n \frac{n^2}{n^3 + 1} \]

OpenStudy (amistre64):

the limit of terms goes to 0 is the sequence decreasing? those are the only criteria i know of to determine convergence with the alternating series thrm

OpenStudy (amistre64):

lucy! you have some splainin' to do ....

OpenStudy (anonymous):

@amistre64 I remember what the choices were now: A) Absolutely converges B) Conditionally converges C) Diverges D) Not enough information E) None of the Above The accepted answer was B, but since this series is just an expression, I really can't see what could be conditional about it. Alas, my dispute was denied, so ¯\_(ツ)_/¯

OpenStudy (amistre64):

you misspelt additionally

OpenStudy (amistre64):

so we know its convergent, the question is to its absoluteness or conditionment ... love making new words

OpenStudy (anonymous):

Heh, indeed I did. The only places I've seen conditional convergence are in functions. e.g.: \[f(x) = \sum_{n=1}^{\infty}x^n\], which of course will only converge if the condition \(|x| < 1\) is met.

OpenStudy (amistre64):

A series is called A.C. if sum|an| is convergent. A series is called C.C. if sum|an| is divergent, but sum an is convergent. so the question is: how does your material define these terms?

OpenStudy (amistre64):

you agreed that sum n^2/(n^3+1) diverges so sum |an| is divergent but we showed that sum an is convergent by series test. so it is CC right?

OpenStudy (amistre64):

'no function variables' is not a condition of CC by the way

OpenStudy (anonymous):

@amistre64 Ah, yes, our text definied it differently. I remember realizing that last year. Our teacher taught us that ac/cc was converge with or without absolute value, but our text said that ac/cc is "converges everywhere"/"converges on an interval".

OpenStudy (amistre64):

you would do well to post the definition your course provides in order for us to assess it better :)

OpenStudy (amistre64):

the actual courses wording; not your .. understanding .. ot it to prevent bias.

OpenStudy (anonymous):

@amistre64 Oh I don't have the text with me anymore. I took BC calc last year, but this year I'm not in any calc class. I do however remember that our teacher told us to ignore its for that chapter. The definitions of AC/CC were along the lines of: \[ f(x) = \sum_{k=k_0}^n a_n\] will absolutely converge iff the sum is finite over \(\Re\), and will conditionally converge iff the sum is finite over a set \(A \subset \Re \). I'll see if any classmates can send me a copy of the definition though.